OpenStudy (anonymous):

what is the distance between the two endpoints (-3,3)(1,-3)?

OpenStudy (johnweldon1993):

Distance formula would work \[\sqrt{(x_2 - x_1)^2 + (y_2 - y_1)^2}\] You have 2 points. (-3,3)(1,-3) (x1,y1)(x2,y2) Plug them in to get your distance

OpenStudy (anonymous):

72.1!!!

OpenStudy (johnweldon1993):

\[\sqrt{(x_2 - x_1)^2 + (y_2 - y_1)^2}\] \[\sqrt{(1 + 3)^2 + (-3 - 3)^2}\] \[\sqrt{(4)^2 + (-6)^2}\] \[\sqrt{16 + 36}\] \[\sqrt{52}\] Is what I get

OpenStudy (anonymous):

ah yes your correct but i was looking for decimal form so we're both right.

OpenStudy (anonymous):

**7.2

OpenStudy (johnweldon1993):

Oh you meant 7.21 ...okay then yes! lol